OpenStudy (anonymous):

Factor the following polynomial completely. 25a4 + 40a2 + 16 =

OpenStudy (shamim):

25a^4=(5a^2)^2

OpenStudy (shamim):

16=4^2

OpenStudy (shamim):

try to use the formula (a+b)^2=a^2+2ab+b^2

OpenStudy (anonymous):

Thanks, that helps a lot

OpenStudy (shamim):

u r most welcome